### v1

# arXiv:2603.00589v1[cs.CV] 28 Feb 2026

Abstract
Visual autoregressive (VAR) models have recently emerged as a promising alternative for image generation, offering stable training, non-iterative inference, and high-fidelity synthesis through next-scale prediction. This encourages the exploration of VAR for image super-resolution (ISR), yet its application remains underexplored and faces two critical challenges: locality-biased attention, which fragments spatial structures, and residual-only supervision, which accumulates errors across scales, severely compromises global consistency of reconstructed images. To address these issues, we propose AlignVAR, a globally consistent visual autoregressive framework tailored for ISR, featuring two key components: (1) Spatial Consistency Autoregression (SCA), which applies an adaptive mask to reweight attention toward structurally correlated regions, thereby mitigating excessive locality and enhancing long-range dependencies; and (2) Hierarchical Consistency Constraint (HCC), which augments residual learning with full reconstruction supervision at each scale, exposing accumulated deviations early and stabilizing the coarse-to-fine refinement process. Extensive experiments demonstrate that AlignVAR consistently enhances structural coherence and perceptual fidelity over existing generative methods, while delivering over 10× faster inference with nearly 50% fewer parameters than leading diffusion-based approaches, establishing a new paradigm for efficient ISR.
1 Introduction
Figure 1: Comparison between the VARSR and AlignVAR. AlignVAR enhances VAR by introducing an adaptive consistency mask for intra-scale modeling and full reconstruction supervision for inter-scale alignment.
Generative models, empowered by strong learned priors, have revolutionized the field of image super-resolution (ISR). Among them, Generative Adversarial Networks (GANs) [17, 19, 25] and Diffusion Models [45, 34, 23] have emerged as the two dominant paradigms. However, each of these approaches suffers from inherent limitations. GAN-based methods [39, 17, 52, 48], while enhancing perceptual realism, often exhibit training instability and tend to introduce visually inconsistent artifacts [43]. Diffusion-based approaches [43, 34, 23, 7, 45] exploit powerful generative priors to achieve high-fidelity reconstruction, yet their iterative denoising process incurs heavy computational costs, substantially limiting their practicality [4].
These challenges have motivated the exploration of alternative generative paradigms for ISR, among which visual autoregressive (VAR) modeling has recently emerged as a promising direction [30, 41]. The coarse-to-fine prediction strategy of VAR naturally aligns with the hierarchical nature of ISR, which progressively restores image details across scales. The recent VARSR [27], as illustrated in Fig. 1, pioneers the application of VAR to ISR, demonstrating its initial feasibility. However, this seminal attempt also reveals a fundamental conflict: the model’s inherent locality bias and the cumulative error propagation across scales jointly undermine the global consistency required for ISR, resulting in degraded global coherence in the reconstructed images.
Figure 2: Comparison of attention distribution. Visualization of attention maps for VARSR and AlignVAR shows that VARSR exhibits highly localized attention concentrated in nearby regions, whereas AlignVAR captures broader contextual dependencies through the proposed Spatial Consistency Autoregression (SCA), thereby enhancing spatial coherence within each scale.
Building on this observation, we conduct a systematic analysis of the VARSR framework and identify two coupled root causes underlying its failure to maintain global consistency. The first is Spatial Inconsistency. The self-attention mechanism in VAR models exhibits a strong locality bias, with attention weights concentrated almost exclusively on adjacent regions, as illustrated in Fig. 2. This restricted receptive field limits the integration of global context, resulting in spatially disjoint artifacts such as fragmented textures and structural distortions, as shown in Fig. 3. The second is Hierarchical Inconsistency, which we analyze by injecting random perturbations into different scales and observing their influence on the final reconstruction, as shown in Fig. 4. These perturbations induce color shifts and structural misalignments, indicating that residual-only supervision allows small prediction errors from coarser scales to propagate and amplify through the hierarchy [15]. We argue that these issues stem from a common origin: the lack of explicit consistency constraints in both the intra-scale (spatial) and inter-scale (hierarchical) dimensions.
To address these limitations, we propose AlignVAR, a visual autoregressive framework designed to achieve globally consistent ISR. As shown in Fig. 1, AlignVAR introduces two complementary mechanisms that collaboratively promote coherence within and across scales by reweighting spatial attention and recalibrating hierarchical dependencies. Specifically, the Spatial Consistency Autoregression (SCA) adaptively reweights attention to emphasize structurally correlated regions rather than local neighborhoods, enabling the model to aggregate long-range context and maintain spatial continuity. Meanwhile, the Hierarchical Consistency Constraint (HCC) recalibrates cross-scale dependencies through a hierarchical supervision objective that enforces both residual and full-scale consistency between predictions and ground-truth representations. By allowing each scale to correct contextual deviations before they propagate, HCC effectively suppresses error accumulation and stabilizes refinement throughout the hierarchy. SCA and HCC form a unified consistency-driven autoregressive formulation that restores fine textures while preserving large-scale structural fidelity, yielding reconstructions that are both perceptually accurate and globally coherent.
Figure 3: Spatial inconsistency results in texture discontinuities, structural distortions.
Figure 4: Hierarchical inconsistency results in color shifts and structural misalignment.
In summary, we make the following key contributions:
  • We revisit visual autoregression for ISR and identify two fundamental sources of inconsistency in existing VAR-based frameworks: spatial inconsistency arising from locality bias and hierarchical inconsistency caused by cumulative error propagation.
  • We propose AlignVAR, a globally consistent VAR framework that enhances spatial coherence and hierarchical alignment through two complementary components: Spatial Consistency Autoregression (SCA) and Hierarchical Consistency Constraint (HCC).
  • Extensive experiments demonstrate that AlignVAR substantially improves reconstruction coherence and perceptual quality, establishing a strong benchmark for autoregressive real-world ISR.
2 Related Works
Image super-resolution.
Early image super-resolution methods learn deterministic mappings from low-resolution to high-resolution images under simplified degradations such as bicubic downsampling, limiting their generalization to real-world scenarios [8, 10, 21, 22, 53, 24]. To overcome this, GAN-based approaches employ adversarial losses to approximate natural image distributions [17, 16, 39, 48, 52]. Methods like BSRGAN [49] and Real-ESRGAN [38] introduce blind degradation modeling for robust real-world recovery, but often suffer from unstable training and unnatural artifacts [43]. Recently, diffusion-based SR methods such as DiffBIR [23], StableSR [34], SeeSR [43], PASD [45], and PiSA-SR [29] leverage generative priors from pretrained diffusion models to enhance perceptual fidelity. Despite their strong visual performance, these models rely on iterative denoising, which incurs high computational cost and may produce hallucinated or inconsistent details [4, 26].
Visual autoregressive models.
Autoregression, motivated by its success in large language models [31, 1], has recently gained traction in visual generation tasks. Most approaches discretize latent features using vector quantizers [32, 18] and predict tokens sequentially [46, 37, 6, 41, 11, 47]. However, next-token prediction over flattened sequences often breaks spatial structure, which makes it difficult to generate coherent high-resolution content [30]. Visual autoregressive modeling [30] addresses this limitation through next-scale prediction, reconstructing images progressively across multiple scales and achieving strong generative performance. This paradigm has also been explored in image restoration [36] and further extended to ISR through VARSR [27], which reconstructs HR images in a coarse-to-fine manner. Nevertheless, VARSR still suffers from locality bias and cumulative error propagation, leading to both spatial and hierarchical inconsistency.
3 Preliminaries
In next-scale visual autoregression, image reconstruction is performed hierarchically within the latent space of a vector-quantized variational autoencoder (VQ-VAE) [32]. At each scale $k$, the latent feature ${f}_{k}$ is computed as the residual between the target latent f{f} and the upsampled reconstructions from all previous coarser scales:
${f}_{k}={f}-\sum_{m=1}^{k-1}\mathrm{upsample}\!\big(\mathrm{lookup}(\mathcal{V},r_{m})\big),$ (1)
where $\mathcal{V}$ denotes the shared codebook, and $\mathrm{lookup}(\mathcal{V},r_{m})$ retrieves the embedded vectors corresponding to tokens $r_{m}$.
Each latent position is quantized to the nearest codebook entry, yielding the discrete token map $r_{k}$:
$r_{k}(i,j)=\arg\min_{v\in[|\mathcal{V}|]}\big\|\mathrm{lookup}(\mathcal{V},v)-{f}_{k}(i,j)\big\|_{2}^{2}.$ (2)
Across scales, the overall autoregressive process follows a coarse-to-fine factorization:
$p(r_{1},r_{2},\dots,r_{K})=\prod_{k=1}^{K}p_{\theta}(r_{k}\mid r_{1},r_{2},\dots,r_{k-1},c),$ (3)
where $c$ represents the conditional latent derived from the low-resolution input.
4 Methods
Figure 5: Overall architecture of the proposed AlignVAR. AlignVAR comprises two complementary components: a Spatial Consistency Autoregression (SCA) that performs scale-wise prediction and reweights intra-scale features using adaptive masks, and a Hierarchical Consistency Constraint (HCC) that jointly supervises residual and full representations to recalibrate inter-scale dependencies.
4.1 Motivation and Empirical Observation
We analyze the spatial behavior of visual autoregression (VAR) in image super-resolution by visualizing its attention distribution. As shown in Fig. 2, VARSR [27] exhibits highly localized attention responses, where most attention mass is confined to narrow regions, revealing a strong local bias. This bias limits long-range contextual aggregation and weakens spatial correlations among distant yet related structures. These observations indicate that sequential tokenization and limited receptive field lead to a non-conservative spatial information flow, resulting in spatial inconsistency within each scale. This motivates our first question:
Q1: How can we mitigate the local bias of VAR to achieve spatial consistency within each scale?
Hierarchical inconsistency caused by error accumulation.
Beyond spatial coherence, the hierarchical prediction process of VAR inherently suffers from cumulative prediction errors, as shown in Fig. 4. In the next-scale prediction paradigm, each latent ${r}_{k}$ is estimated based on imperfect outputs from preceding scales ${r}_{1:k-1}$. Any deviation in coarse-scale predictions shifts the conditional distribution $p({r}_{k}\mid{r}_{1:k-1})$, leading to a compounding effect in which errors are propagated and amplified across scales. We attribute this problem to residual-only supervision used in existing VAR frameworks, which leaves intermediate latent representations under-constrained and destabilizes the generative hierarchy. This motivates our second question:
Q2: How can we recalibrate intermediate predictions to maintain hierarchical consistency across scales?
4.2 Overview
Motivated by the two issues identified above, we propose AlignVAR, a globally consistent visual autoregressive framework for ISR. The overall architecture of AlignVAR is illustrated in Fig. 5. Briefly, AlignVAR is composed of two primary modules: Spatial Consistency Autoregression (SCA) and Hierarchical Consistency Constraint (HCC). The SCA expands the effective receptive field and applies a learnable structure-aware mask to mitigate local attention bias (§4.3), while the HCC enforces full-scale latent alignment to recalibrate intermediate predictions and suppress cumulative drift across scales (§4.4). Together, these designs enable AlignVAR to produce perceptually faithful and structurally coherent high-resolution reconstructions.
4.3 Spatial Consistency Autoregression (SCA)
The core idea of SCA is to enhance spatial coherence within the autoregressive process. Rather than relying on order-based attention with an inherently local bias, SCA introduces a structure-aware conditioning principle that reweights contextual dependencies.
Specifically, SCA preserves the coarse-to-fine autoregressive structure but replaces the plain context $(r_{1},\dots,r_{k-1})$ with a structure-aware counterpart. Let $\tilde{r}_{k}$ denote the reweighted tokens at scale $k$. The joint distribution under SCA becomes
$p_{\text{SCA}}(\tilde{r}_{1},\dots,\tilde{r}_{K})=\prod_{k=1}^{K}p_{\theta}\big(\tilde{r}_{k}\mid\tilde{r}_{1},\tilde{r}_{2},\dots,\tilde{r}_{k-1},c\big),$ (4)
where the next-scale prediction at level $k$ is conditioned on structure-aligned representations of all coarser scales. SCA thus remains defined only across scales, while spatial consistency of the context is enforced through structure-aware reweighted tokens $\tilde{r}_{1:k-1}$.
Structure-aware reweighting field.
As shown in the bottom middle part of Fig. 5, we use structure-aware guidance extracted from the low-resolution input $I_{LR}$ to capture structural cues corresponding to potential edge or texture regions:
$s=\big|\mathrm{Laplacian}(I_{LR})\big|,\qquad\bar{s}_{k}=\mathrm{norm}\!\big(\mathrm{Down}_{k}(s)\big),$ (5)
where $\mathrm{Laplacian}$ represents a Laplacian operator [5], chosen for its sensitivity to second-order structural changes, enabling the guidance map to highlight edges for effective spatial reweighting. $\mathrm{Down}_{k}(\cdot)$ denotes scale-specific downsampling that matches the spatial resolution of the $k$-th latent map, and $\mathrm{norm}(\cdot)$ rescales values to $[0,1]$.
By normalizing across scales, the structural guidance $\bar{s}_{k}$ remains comparable in magnitude and semantics, allowing the model to leverage multi-scale geometric consistency rather than relying solely on local correlations.
Subsequently, we use a lightweight MLP-based mask generator $\mathcal{M}_{\phi}$ to predict a spatial modulation field $m_{k}$ from the autoregressive tokens $r_{k}$ and the corresponding structural guidance $\bar{s}_{k}$:
$m_{k}=\sigma\!\big(\mathcal{M}_{\phi}([\,r_{k},\,\bar{s}_{k}\,])\big),$ (6)
where $\cdot\,$] denotes channel-wise concatenation and $\sigma$ is a sigmoid function. This modulation field $m_{k}$ serves as a structure-aware reweighting map that adaptively adjusts the token responses according to the reliability of the underlying structure. Regions with clear geometric cues are assigned higher weights, encouraging the model to attend to stable, well-defined features, while uncertain or textureless areas are softly suppressed to reduce local noise amplification. Finally, the structure-aware reweighted tokens are produced through a learnable spatial field via token gating:
$\tilde{r}_{k}=(1+m_{k})\odot r_{k},$ (7)
with $\odot$ denoting element-wise multiplication. This operation can be viewed as a spatially adaptive gain control that modulates the strength of autoregressive context propagation. By reinforcing structure-aligned activations and suppressing unstable ones, the reweighting field guides the model to preferentially propagate information along reliable structural paths. This selectively amplifies long-range correlations between semantically related regions, effectively mitigating local bias and enabling broader contextual aggregation within each scale.
While SCA enforces spatial coherence within each scale, the hierarchical dependencies across scales remain unregularized. Section 4.4 introduces the Hierarchical Consistency Constraint to align inter-scale representations and suppress error propagation through the hierarchy.
4.4 Hierarchical Consistency Constraint (HCC)
Figure 6: Comparison of multi-scale reconstructions. AlignVAR alleviates cumulative error propagation across scales and preserves more consistent structural and textural details than VARSR [27].
In the original next-scale autoregressive formulation, the Cross Entropy (CE) loss [30] supervises only the residual tokens predicted at each scale. While effective for local refinement, this residual-only supervision cannot correct errors accumulated from coarser levels, allowing scale-wise biases to propagate upward. To address this issue, we introduce HCC, which provides full-scale latent supervision to explicitly recalibrate hierarchical dependencies. By aligning predicted latent representations with their complete multi-scale ground truths, HCC enforces global consistency across the entire autoregressive hierarchy.
Full-scale representation.
Given a high-resolution image ${I}_{HR}\in\mathbb{R}^{C\times H_{K}\times W_{K}}$, the VAE encoder $\mathcal{E}$ produces its latent feature representation
${z}=\mathcal{E}({I}_{HR})\in\mathbb{R}^{C\times H_{K}\times W_{K}}.$ (8)
To construct scale-specific ground truths for hierarchical supervision, we spatially downsample z{z} to each target resolution $S_{k}$ and quantize the result into discrete tokens:
${u}_{\text{gt}}^{k}=\mathcal{Q}\!\big(\mathrm{Down}({z},S_{k})\big),\qquad S_{k}\in\{S_{1},S_{2},\dots,S_{K}\},$ (9)
where $\mathrm{Down}(\cdot)$ denotes spatial downsampling to the resolution of scale $k$, and $\mathcal{Q}(\cdot)$ denotes the shared vector quantizer. Unlike the residual tokens used in the standard CE loss, ${u}_{\text{gt}}^{k}$ represents the complete latent state at each scale, thus providing a stronger supervision target.
HCC loss.
During autoregressive reconstruction, the model predicts residual tokens $\hat{{r}}_{\text{pred}}^{k}$ for each scale. The corresponding cumulative (full-scale) latent prediction is obtained by aggregating residuals from all previous scales:
$\hat{{u}}_{\text{pred}}^{k}=\hat{{u}}_{\text{pred}}^{k-1}+\hat{{r}}_{\text{pred}}^{k},\qquad\hat{{u}}_{\text{pred}}^{(0)}=\mathbf{0}.$ (10)
We design the HCC loss to impose a multi-scale hierarchical alignment constraint between these cumulative predictions and the corresponding ground-truth latents:
$\mathcal{L}_{\text{HCC}}=\sum_{k=1}^{K}\big\|\hat{{u}}_{\text{pred}}^{k}-{u}_{\text{gt}}^{k}\big\|_{2}^{2}.$ (11)
This inter-scale calibration prevents error propagation and aligns the semantic content across the hierarchy.
By extending the supervision from residuals to full latent representations, the HCC loss bridges the semantic gap between hierarchical scales and stabilizes the coarse-to-fine refinement process. As shown in Fig. 6, AlignVAR reconstructs finer and more consistent textures at higher scales compared to VARSR [27], indicating that the proposed recalibration effectively alleviates the accumulation and propagation of local errors. Together with SCA’s spatial regularization, it establishes a unified mechanism that preserves intra-scale coherence and inter-scale consistency throughout the autoregressive super-resolution process.
4.5 Training Objective
Our training process is performed under the teacher-forcing paradigm [30], where the prediction at each scale is conditioned on the reweighted ground-truth tokens $\tilde{r}_{\mathrm{gt}}^{1:k-1}$ and the low-resolution latent feature $c$, providing a stable and reliable supervision path throughout the entire multi-scale prediction procedure. The autoregressive predictor models a categorical distribution over the codebook entries at each spatial position $(i,j)$, optimized via a standard CE objective commonly used in modern generative modeling:
$\mathcal{L}_{\mathrm{CE}}=-\sum_{k=1}^{K}\sum_{(i,j)}\log p_{\theta}\big(r_{\mathrm{gt}}^{k}(i,j)\mid\tilde{r}_{\mathrm{gt}}^{1:k-1},c\big).$ (12)
Finally, the overall objective of AlignVAR combines both intra-scale and inter-scale learning signals, ensuring that each component contributes meaningfully to the reconstruction quality:
$\mathcal{L}_{\mathrm{total}}=\mathcal{L}_{\mathrm{CE}}+\lambda\mathcal{L}_{\mathrm{HCC}},$ (13)
where $\lambda$ is a hyperparameter that balances the two components. Joint optimization of $\mathcal{M}_{\phi}$ and $p_{\theta}$ encourages spatially coherent predictions at each scale and hierarchically consistent dependencies across the reconstruction process, ultimately improving both stability and performance.
Table 1: Comparison with state-of-the-art methods on synthetic and real-world benchmarks. The best and second-best results are highlighted in bold red and underline blue, respectively.
Datasets Metrics GAN-based Diffusion-based VAR-based
BSRGAN Real-ESR SwinIR LDM StableSR DiffBIR PASD UPSR VARSR AlignVAR
[…]
5 Experimental Results
5.1 Experimental Setups
Datasets.
AlignVAR is trained on LSDIR [20] and the first 10K face images from FFHQ [13], where low-resolution counterparts are synthesized using the Real-ESRGAN degradation pipeline [38]. For evaluation, we adopt both synthetic and real-world benchmarks. The synthetic validation set (DIV2K-Val) is built by randomly cropping 3K patches from DIV2K [3], while real-world evaluation is conducted on DRealSR [42] and RealSR [33]. Following [30], all HR images are standardized to $512\times 512$ and their LR inputs to $128\times 128$.
Implementation details.
AlignVAR is trained using a scale-wise autoregressive transformer backbone consisting of 24 transformer blocks. The model is initialized from pretrained VAR weights [30] for faster convergence. Training uses AdamW [2] with a batch size of 32, weight decay of $5\times 10^{-2}$, and an initial learning rate of $5\times 10^{-5}$ decayed via cosine annealing. The model is optimized for 100 epochs, with the loss balancing coefficient $\lambda$ in Eq. 13 set to 1.0. All experiments are conducted on 8 NVIDIA H100 GPUs.
Figure 7: Qualitative comparisons with representative methods.
Evaluation metrics.
To comprehensively evaluate both fidelity and perceptual quality, we adopt a diverse set of reference-based and no-reference metrics. For fidelity, PSNR and SSIM [40] are computed on the Y channel of the YCbCr color space. Perceptual similarity is measured using LPIPS [51] and DISTS [9], while distributional alignment is assessed with FID [12]. We further employ no-reference metrics, including MANIQA [44], MUSIQ [14], and CLIPIQA [35], to evaluate visual naturalness.
Compared models.
We compare AlignVAR with state-of-the-art (SOTA) methods in two categories. The first includes GAN-based methods: BSRGAN [49], Real-ESRGAN [38], and SwinIR-GAN [21]. The second covers diffusion-based methods: LDM [28], StableSR [34], DiffBIR [23], PASD [45], and UPSR [50]. We also include the recent autoregressive method VARSR [27], retrained on our dataset for fairness. Other competing models are evaluated using official implementations and pretrained weights.
5.2 Comparison with SOTA
Quantitative comparisons.
As shown in Table 1, AlignVAR performs strongly across both synthetic and real-world benchmarks, substantially improving overall perceptual quality. On the DIV2K-Val dataset, it outperforms GAN-based and diffusion-based methods in all perceptual metrics, achieving the lowest FID of 25.71 and the best LPIPS of 0.2955 among all compared models. On RealSR, AlignVAR significantly increases MUSIQ from 66.65 to 68.53 and CLIPIQA from 0.5953 to 0.6784 compared to VARSR [27], demonstrating its consistent effectiveness. Although AlignVAR does not achieve the highest fidelity metrics, this is expected because fine details in LR images are severely degraded and cannot be fully recovered by any existing method. AlignVAR consistently produces high-quality super-resolution results that better align with human visual perception, offering perceptually realistic, visually coherent, and structurally consistent reconstructions.
Qualitative comparisons.
Fig. 7 presents visual comparisons with representative ISR methods. Some artifacts highlight common shortcomings of existing paradigms: GAN-based models often cause local distortions and jagged edges, while diffusion-based models may hallucinate textures and weaken structural alignment. In contrast, AlignVAR reconstructs sharp edges, coherent textures, and natural color transitions, better aligning with human visual perception. Notably, when true high-frequency details are irrecoverably lost in the low-resolution input, AlignVAR preserves the recoverable structures and generates perceptually plausible textures. This explains its clearly superior no-reference metric scores, even if the reference metric scores in Table 1 are not correspondingly higher.
Table 2: Complexity comparison.
Models Params Steps Inference Time
StableSR 1409.1M 200 15.32s
DiffBIR 1900.4M 20 5.03s
PASD 1716.7M 50 5.94s
UPSR 1530.9M 5 2.79s
VARSR 1102.9M 10 0.52s
AlignVAR 1056.5M 10 0.43s
Table 3: Ablation study of the SCA. “w/o” denotes removing SCA, “RI” indicates random input, and “SG” represents the structural guidance used in our model.
Metrics RealSR DrealSR
w/o SCA RI SG w/o SCA RI SG
[…]
Complexity comparisons.
As shown in Table 2, AlignVAR achieves significantly higher efficiency than diffusion-based models and VARSR [27]. Diffusion-based methods require iterative refinement, leading to long inference times even with a limited number of sampling steps. In contrast, AlignVAR reconstructs a $512\times 512$ image in only 0.43 seconds, which is over $10\times$ faster than PASD and more than $5\times$ faster than the 5-step UPSR. This efficiency stems from the scale-wise prediction strategy, where early scales involve fewer tokens and thus incur lower computational cost. Compared with the autoregressive baseline VARSR, which incorporates an additional diffusion refiner at the expense of increased parameters and latency, AlignVAR removes this component and adopts a lightweight mask generator that introduces negligible overhead. Thus, AlignVAR achieves better overall performance with fewer parameters and faster inference. Further analysis of the efficiency advantage of AlignVAR is provided in the Supplementary Material.
Figure 8: Comparison of supervision domains for HCC.
5.3 Ablation Study and Analysis
Effectiveness of SCA.
We evaluate the effectiveness of SCA on the RealSR and DRealSR datasets, as shown in Table 3. Three configurations are compared: removing SCA, using random input for the mask generator (RI), and applying the proposed structural guidance (SG). Removing SCA slightly improves fidelity metrics but noticeably degrades perceptual scores, while random input leads to unstable textures and weakened visual coherence. These results demonstrate that SCA effectively exploits structural guidance to enhance spatial consistency and achieves a better balance between fidelity and perceptual quality.
Effectiveness of HCC.
We evaluate the impact of HCC on the RealSR and DRealSR datasets, as shown in Table 4. Introducing HCC consistently improves both fidelity and perceptual metrics, indicating that aligning full-scale latent representations strengthens cross-scale information flow and stabilizes hierarchical prediction. We further compare applying this constraint in latent space versus in pixel space and observe that latent space supervision achieves noticeably better perceptual coherence, as illustrated in Fig. 8. Overall, HCC effectively suppresses accumulated errors and enhances inter-scale consistency, yielding reconstructions that are both more accurate and visually coherent.
Influence of the balancing coefficient.
We study the effect of the balancing coefficient $\lambda$ in Eq. 13 on the RealSR dataset, as shown in Table 5. When $\lambda$ increases from 0.5 to 1.0, both perceptual and fidelity metrics improve, with the best perceptual quality achieved at $\lambda=1.0$. Further increasing $\lambda$ beyond 1.0 slightly enhances distortion-oriented scores but leads to a decrease in perceptual quality. These results suggest that $\lambda=1.0$ provides better overall performance in terms of perceptual quality while maintaining competitive fidelity scores.
Analysis of the Adaptive Mask.
Fig. 9 visualizes the structural guidance, the corresponding initial and learned masks, and the reweighted feature map. The structural guidance captures coarse edges and contours, which are transferred to the initial mask as a spatial bias. After training, the learned mask becomes sharper and more selective, concentrating on semantic boundaries such as windmill blades and Chinese characters while suppressing activations in smooth background regions. The reweighted feature map highlights these structural areas with clearer and more coherent responses, suggesting that the mask adaptively emphasizes perceptually important regions. Overall, this process enables more structure-aware feature interaction and improves spatial consistency in the reconstructed images.
Table 4: Ablation on the Hierarchical Consistency Constraint.
Metrics RealSR DrealSR
w/o HCC HCC w/o HCC HCC
[…]
Table 5: Effect of the balancing coefficient $\lambda$ on RealSR.
[…]
$\lambda=0.5$ 25.83 0.2874 0.2184 0.4403 67.56
$\lambda=1.0$ 26.11 0.2871 0.2123 0.4553 68.53
$\lambda=1.5$ 26.21 0.2785 0.2107 0.4465 67.85
$\lambda=2.0$ 26.35 0.2703 0.2079 0.4337 67.03
Analysis of HCC.
To further understand the effect of HCC, we analyze its impact on hierarchical recalibration and robustness to perturbations. Fig. 10 (left) plots the mean squared error (MSE) at each prediction scale. AlignVAR without HCC already reduces overall errors compared with VARSR, and adding HCC further lowers the MSE at early and middle scales, indicating that HCC serves as a coarse-to-fine recalibration mechanism that corrects global structures before finer refinement. For robustness evaluation, we inject identical random perturbations into latent features at scales 2, 5, and 8 and report MUSIQ scores in Fig. 10 (right). AlignVAR shows the smallest MUSIQ degradation under all noise levels, confirming that HCC improves inter-scale stability against cumulative perturbations.
Figure 9: Analysis of the adaptive mask at different training stages. SG denotes the structural guidance extracted from the LR input.
Figure 10: Quantitative analysis of the Hierarchical Consistency Constraint on RealSR. Shown are the mean prediction errors across scales (left) and the MUSIQ scores under identical random perturbations injected at different scales (right).
6 Conclusion
We present AlignVAR, a globally consistent visual autoregressive framework for image super-resolution. Through a detailed and comprehensive technical analysis of inconsistency issues in existing VAR models, we identify spatial local bias and hierarchical error propagation as the main bottlenecks severely hindering coherent and accurate reconstruction. To alleviate these issues, we introduce Spatial Consistency Autoregression to adaptively reweight fine-grained intra-scale attention and Hierarchical Consistency Constraint to recalibrate multi-level inter-scale dependencies through joint residual and full-scale supervision. Extensive experiments on both synthetic and real-world benchmarks consistently demonstrate that AlignVAR achieves an excellent overall balance between fidelity and perceptual quality. This work provides a new perspective for achieving globally coherent and reliable visual autoregression for high-quality image super-resolution.